find the third of four consecutive odd integers, given that the sum of the smallest two integers is four times the largest
Answer by ankor@dixie-net.com(22740)

find the third of four consecutive odd integers, given that the sum of the smallest two integers is four times the largest
:
Let = x the third of the 4 consecutive odd integers
:
The 4 integers:
(x-4), (x-2), x, (x+2)
:
"the sum of the smallest two integers is four times the largest"
:
(x-4) + (x- 2) = 4(x+2)
:
x + x - 4 -2 = 4x + 8
:
2x - 6 = 4x + 8
:
2x - 4x = 8 + 6
:
-2x = 14
:
x = 14/-2
:
x = -7 is the 3rd integer
:
:
Check, the 4 integers, -11, -9, -7, -5
-11 + (-9) = 4(-5)
-20 = -20